OpenStudy (anonymous):

Little help? state the amplitude period and transformation of the function y=2-3sin(pi/3x-pi)

OpenStudy (kropot72):

\[y=2-3\sin (\frac{\pi}{3}x-\pi)\] Is that it?

OpenStudy (anonymous):

yes it is

OpenStudy (kropot72):

The general equation is: \[y=A \sin (Bx+C)+D\] Amplitude = |A| \[Period\ T=\frac{2 \pi}{B}\] \[Phase\ shift=\frac{C}{B}\] \['y'\ shift=D\] Using that, can you identify the amplitude?

OpenStudy (anonymous):

ok is the amp -3?

OpenStudy (kropot72):

It might help if I rearrange the given equation as follows: \[y=-3\sin (\frac{\pi}{3}x-\pi)+2\]

OpenStudy (kropot72):

The amplitude is |-3| = ?

OpenStudy (anonymous):

3 got it thanks!!

OpenStudy (anonymous):

so the period is 2pi/pi/3?

OpenStudy (kropot72):

Good work! Can you simplify it?

OpenStudy (anonymous):

so its 6 and the horizontal shift is pi?

OpenStudy (kropot72):

The period is 6. The phase shift is: \[\frac{- \pi}{\frac{\pi}{3}}\]

OpenStudy (anonymous):

ok thanks got it!

OpenStudy (kropot72):

Good work! You're welcome :)